OpenStudy (anonymous):

what is the first 6 terms in the sequence 12-2n

OpenStudy (anonymous):

and the 20th term

OpenStudy (midhun.madhu1987):

Put the values of n=1,2,3...6

OpenStudy (anonymous):

what is the sequence for 12-2n

OpenStudy (anonymous):

but thanks for helping

OpenStudy (midhun.madhu1987):

12 - 2*1 = 12-2 = 10 12 - 2*2 = 12 - 4 = 8 ....

OpenStudy (anonymous):

thank you very much that really helped